### v1

# Quaternionic Pole Placement via Companion Forms and the Ackermann Formula

Abstract
We present an extension of state-feedback pole placement for quaternionic systems, based on companion forms and the Ackermann formula. For controllable single-input quaternionic LTI models, we define a companion polynomial that right-annihilates its companion matrix, characterize spectra via right-eigenvalue similarity classes, and prove coefficient-matching design in controllable coordinates. We then derive a coordinate-free Ackermann gain expression valid for real target polynomials, and state its scope and limitations. Short examples demonstrate correctness, practical use, and numerical simplicity.
I Introduction
Quaternions provide a natural extension of real and complex numbers: like $\mathbb{R}$ and $\mathbb{C}$, $\mathbb{H}$ is a division algebra, but—crucially—multiplication is noncommutative. This changes some rules in systems and control. Eigenvalues must be interpreted carefully (right vs. left) and—on the right—they are defined only up to quaternionic similarity classes. Determinants and a full Cayley–Hamilton theorem do not carry over in the usual way, and even polynomial evaluation is order-dependent, so right and left zeros need not coincide. As a result, several classical tools are not yet fully developed—or are simply missing—when one works directly over $\mathbb{H}$; see, e.g., foundational results on quaternionic LTI stability [19] and linear algebra [21].
Quaternion models are a natural choice whenever states or operators are genuinely hypercomplex. Beyond rigid-body attitude kinematics, representative areas include: spin-$\tfrac{1}{2}$ quantum dynamics [25] and qubit gate control—e.g., flatness-based single-qubit gates [5]; NMR/MRI pulse design—e.g., quaternion descriptions of selective-pulse propagators [6]; Polarization/bivariate signal processing and colour imaging—e.g., Quaternion-MUSIC for DOA–polarization [16], quaternion Fourier transforms for colour images [22], and a recent overview [13]; and learning and computation with hypercomplex variables—e.g., quaternion neural controllers [23], quaternion CNNs for speech [18], and foundational links to quantum computing [15].
Other interpretations of quaternions are already used in control of special classes of systems—for example, unit quaternions as rotation parameters, complex $2{\times}2$ representations, or real four-vector embeddings. In this note, we adopt the hypercomplex algebraic viewpoint, where signals, states, and operators naturally reside in $\mathbb{H}$. This keeps the calculus aligned with the modeling (no repeated conversions to $\mathbb{R}$ or $\mathbb{C}$), makes spectral notions explicit (right eigenvalues and their similarity classes), and meshes naturally with companion constructions used in design.
The primary contribution of this work is to lift standard real/complex state-space methods to quaternionic systems. We tackle a basic, practical task: state-feedback pole placement for single-input, controllable LTI models over $\mathbb{H}$. The results rest on two pillars. First, we develop quaternionic companion forms together with a quaternionic companion polynomial that right-annihilates its companion matrix—playing the role of characteristic/minimal polynomials over fields—and thereby provide a clean, determinant-free framework in controllable (companion) coordinates. Second, we give two equivalent design routes: (i) coefficient matching in the controllable companion form, and (ii) a coordinate-free Ackermann-type expression based only on the controllability matrix and a desired companion polynomial. We also specify the exact scope: the Ackermann route applies when the target polynomial has real (central) coefficients, which suffices to prescribe any combination of real classes and complex-conjugate classes of right eigenvalues.
These workflows resemble their $\mathbb{R}/\mathbb{C}$ counterparts—companion form, one-shot gain, and an Ackermann-style formula—but they differ in substance: over $\mathbb{H}$, right eigenvalues (the poles) are similarity classes (not points), polynomial evaluation is side/ordering-sensitive, and the Ackermann step is valid only for real (central) coefficients. Our design preserves the familiar flow while making these quaternionic nuances explicit.
The note is organized as follows. Section II reviews quaternion algebra and matrix preliminaries; Section III states the quaternionic LTI model. Section IV introduces the controllable companion form and its companion polynomial. Section V formulates pole placement over $\mathbb{H}$. Section VI presents coefficient matching in controllable coordinates; Section VII derives the quaternionic Ackermann formula and delineates its scope; and Section VIII concludes.
II Quaternions and Matrices
Quaternions
A quaternion is a hypercomplex scalar $q=a+b\mathbf{i}+c\mathbf{j}+d\mathbf{k}$ with $a,b,c,d\in\mathbb{R}$, and imaginary units $\mathbf{i},\mathbf{j},\mathbf{k}$ satisfying $\mathbf{i}^{2}=\mathbf{j}^{2}=\mathbf{k}^{2}=\mathbf{i}\mathbf{j}\mathbf{k}=-1$ (hence $\mathbf{i}\mathbf{j}=\mathbf{k}=-\mathbf{j}\mathbf{i},$ $\mathbf{j}\mathbf{k}=\mathbf{i}=-\mathbf{k}\mathbf{j},$ $\mathbf{k}\mathbf{i}=\mathbf{j}=-\mathbf{i}\mathbf{k}$). The set of all quaternions, denoted $\mathbb{H},$ is a 4-dimensional associative but noncommutative division algebra over $\mathbb{R}.$ For $q\in\mathbb{H},$ denote its real and imaginary parts by $\operatorname{Re}(q)=a$ and $\operatorname{Im}(q)=b\mathbf{i}+c\mathbf{j}+d\mathbf{k}$, respectively; the conjugate is $\bar{q}=a-b\mathbf{i}-c\mathbf{j}-d\mathbf{k}$ and the norm is $|q|=\sqrt{q\bar{q}}.$ Throughout we use the scalar (hypercomplex) model of $\mathbb{H}$; see [21] for a more detailed introduction.
Similarity and classes
Two quaternions $q,r$ are similar, written $q\sim r,$ if $q=\alpha^{-1}r\alpha$ for some nonzero $\alpha\in\mathbb{H}.$ Similarity preserves $\operatorname{Re}(\cdot),$ $\left|\cdot\right|,$ and $\left|\operatorname{Im}(\cdot)\right|$ and partitions $\mathbb{H}$ into equivalence classes $\left[\cdot\right]$: real quaternions are isolated classes $\left[a\right]=\{a\}$; If $\operatorname{Im}(q)\neq 0,$ then $[q]$ is a 2-sphere $\{\mathrm{Re}(q)+u\,\left|\operatorname{Im}(q)\right|:u\in\mathrm{span}\{\mathbf{i},\mathbf{j},\mathbf{k}\},\left|u\right|=1\}$ that meets the standard complex line $\{a+b\mathbf{i}:\,a,b\in\mathbb{R}\}$ in exactly two points $\operatorname{Re}(q)\pm i\,\left|\operatorname{Im}(q)\right|.$ For example, $\mathbf{k}]$ contains all imaginary units, including $\pm\mathbf{i}.$
Quaternionic matrices
For $m,n\in\mathbb{N}$, let $\mathbb{H}^{m\times n}$ denote the set of $m\times n$ matrices with entries in $\mathbb{H}.$ We view column vectors as a right $\mathbb{H}$-module and use the standard conjugate transpose $A^{*}=\overline{A}^{\mathsf{T}}.$ Noncommutativity shows up strongly: entrywise conjugation and transpose do not distribute over products in the usual way: in general $\overline{AB}\neq\bar{A}\bar{B}$ and $(AB)^{\mathsf{T}}\neq B^{\mathsf{T}}A^{\mathsf{T}}$. By contrast, the mixed identity holds: $(AB)^{\ast}=B^{\ast}A^{\ast}.$ Likewise, $(A^{\top})^{-1}=(A^{-1})^{\top}$ need not hold over $\mathbb{H}$ (as transpose may destroy invertibility), whereas $(A^{*})^{-1}=(A^{-1})^{*}$ does hold. See e.g. [21] for further subtleties.
Rank and invertibility
For $A\in\mathbb{H}^{m\times n}$, the quaternionic rank, denoted $\operatorname{rank}_{\mathbb{H}}A$, is the maximal number of right–independent columns of $A$ (i.e., if $\sum_{j}a_{j}c_{j}=0$ with coefficients $c_{j}\in\mathbb{H}$ on the right, then all $c_{j}=0$). Equivalently, it is the maximal number of left–independent rows. A square matrix $A\in\mathbb{H}^{n\times n}$ is invertible iff $\operatorname{rank}_{\mathbb{H}}A=n,$ see [21].
Right eigenvalues and spectrum
For $A\in\mathbb{H}^{n\times n}$, a right eigenpair is $(v,\lambda)$ with $v\neq 0$ and
$Av$ $=$ $v\,\lambda,\qquad v\in\mathbb{H}^{n},\ \lambda\in\mathbb{H}.$ (1)
Scaling $v$ on the right by $\alpha\neq 0\in\mathbb{H}$ replaces $\lambda$ by $\alpha^{-1}\lambda\alpha.$ So, right eigenvalues are defined up to quaternionic similarity. Therefore, the right spectrum is the set of similarity classes
$\sigma_{\mathscr{r}}(A):=\{[\lambda]:\exists v\neq 0\text{ such that }Av=v\lambda\}\subset\mathbb{H}/\sim,$
which is invariant under matrix similarity: $A\mapsto S^{-1}AS$ [21]. When a scalar is needed, we choose the unique complex representative in each class with a nonnegative imaginary part and call it the standard eigenvalue. (Left eigenvalues, defined by $Av=\mu v$, are not preserved by matrix similarity and will not be used unless explicitly stated.)
Polynomials and evaluations
Let $\mathbb{H}[\lambda]$ denote the ring of quaternionic polynomials $p(\lambda)=\sum_{k=0}^{n}p_{k}\lambda^{k}$ with $p_{k}\in\mathbb{H}$, where the indeterminate $\lambda$ commutes with all coefficients. Evaluation at a quaternion $q\in\mathbb{H}$, however, is order–dependent: $p(q)_{\mathscr{r}}:=\sum_{k=0}^{n}p_{k}q^{k}$ is the right value and $p(q)_{\ell}:=\sum_{k=0}^{n}q^{k}p_{k}$ is the left value.
Left and right zeros
A right zero of $p$ is $q\in\mathbb{H}$ such that $p(q)_{\mathscr{r}}=0;$ a left zero satisfies $p(q)_{\ell}=0$. Left and right zeros differ only by similarity. If $p(q)_{\ell}=0$, then there is some $\alpha\neq 0$ with $p(\alpha^{-1}q\alpha)_{\mathscr{r}}=0$, and conversely [9]. If all $p_{k}\in\mathbb{R}$ (central coefficients), then $p(q)_{\mathscr{r}}=p(q)_{\ell}$ for all $q$ and the left zeros equal the right zeros.
Isolated and spherical zeros
A quaternionic polynomial has either $n$ (not necessarily distinct) right zeros or infinitely many. A nonreal zero is either isolated (a single point of its class) or spherical (the entire class $[q]$, a 2-sphere). The spherical case occurs exactly when the real quadratic $\chi_{q}(\lambda)=\lambda^{2}-2\,\mathrm{Re}(q)\,\lambda+|q|^{2}$ divides $p(\lambda);$ Real zeros are always isolated since $r]=\{r\}.$ With real coefficients, all zeros are either real (hence isolated) or spherical [20].
Notation
We view column vectors as a right $\mathbb{H}$-module. Let $e_{i}$ denote the $i$th canonical basis vector in $\mathbb{R}^{n}\subset\mathbb{H}^{n}$, i.e., the column with a $1$ in position $i$ and zeros elsewhere; in particular $e_{n}=[0,\ldots,0,1]^{\top}$. Since these entries are real (central), right/left multiplication with scalars is unambiguous when $e_{i}$ appears.
III Quaternionic state–space systems
We study continuous-time LTI state–space models over $\mathbb{H}$
$\dot{x}(t)$ $=$ $Ax(t)+Bu(t),$ (2)
and the output equation plays no role in what follows. For clarity, we restrict to the single-input case: $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$, with state $x(t)\in\mathbb{H}^{n}$ and input $u(t)\in\mathbb{H}$. (When desired, an output $y(t)=Cx(t)+Du(t)$ with $C\in\mathbb{H}^{p\times n}$ and $D\in\mathbb{H}^{p}$ can be appended without affecting the development.)
Stability
Asymptotic (internal) stability is defined exactly as in the real/complex case: with $u\equiv 0$, every trajectory satisfies $x(t)\to 0$ for all $x(0)\in\mathbb{H}^{n}$. Equivalently, (2) is asymptotically stable iff every right–eigenvalue class of AA meets $\mathbb{C}$ in a point $\lambda$ with $\operatorname{Re}\,\lambda<0$; see [19].
Controllability
As in the real/complex case, controllability means the ability to steer the state from any $x_{0}\in\mathbb{H}^{n}$ to any $x_{f}\in\mathbb{H}^{n}$ in finite time by a suitable input. In the single–input setting, the controllability matrix is
$\mathcal{C}$ $=$ $\bigl[\,B\;\;AB\;\;\cdots\;\;A^{n-1}B\,\bigr]\in\mathbb{H}^{n\times n},$ (3)
and the pair $(A,B)$ is completely controllable iff $\mathcal{C}$ is invertible over $\mathbb{H}$ (equivalently, iff $\operatorname{rank}_{\mathbb{H}}\mathcal{C}=n$) [10].
IV Companion Form and Companion Polynomial
In classical (real/complex) linear systems, two key tools are the controllable companion form and the characteristic or minimal polynomial. For quaternionic systems, the controllable companion form still applies. However, characteristic and minimal polynomials do not carry over directly, because there is no simple determinant on quaternions. Instead, we can work with a quaternionic companion polynomial.
Definition 1 (Controllable companion form and companion polynomial).
Let $A_{c}\in\mathbb{H}^{n\times n}$ and $B_{c}\in\mathbb{H}^{n}$ be
$A_{c}$ $=$ $\begin{bmatrix}0&1&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&1\\ -a_{0}&-a_{1}&\ldots&-a_{n-1}\\ \end{bmatrix}$,\;\; B_c = [0 ⋮0 1 ]. (4)
with $a_{0},\ldots,a_{n-1}\in\mathbb{H}$. We say that $(A_{c},B_{c})$ is in (lower) controllable companion (canonical) form. The companion polynomial associated with $(A_{c},B_{c})$ (or with $A_{c}$) is the monic quaternionic polynomial
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{H}[\lambda].$ (5)
If $A\in\mathbb{H}^{n\times n}$ is cyclic—i.e., there exists $B\in\mathbb{H}^{n}$ with $\mathcal{C}$ invertible—then $A$ is similar to the lower controllable companion form. The next theorem makes this precise and gives the construction.
Theorem 1 (Determinant-free companion form for a controllable pair).
Let $A\in\mathbb{H}^{n\times n}$ and $B\in\mathbb{H}^{n}$ be given, and assume $\mathcal{C}=[B,AB,\ldots,A^{n-1}B]$ is invertible over $\mathbb{H}$.
  1. 1.
    Then there exists an invertible similarity $T\in\mathbb{H}^{n\times n}$ such that
    $A_{c}=T^{-1}AT\quad\textrm{ and }\quad B_{c}$ $=$ $T^{-1}B$ (6)
    and the pair $(A_{c},B_{c})$ is in lower controllable companion form (4), where the coefficients $a_{0},\dots,a_{n-1}\in\mathbb{H}$ are uniquely defined by the (right) expansion
    $A^{n}B\;=\;-\sum_{\mathbf{k}=0}^{n-1}A^{k}B\,a_{k}.$ (7)
  2. 2.
    The matrix $T$ is explicitly given by
    $T^{-1}$ $=$ $\begin{bmatrix}t\\ tA\\ \vdots\\ tA^{n-1}\end{bmatrix}$. (8)
    where the first row reads
    $t$ $=$ $[0,0,\ldots,1]\,\mathcal{C}^{-1}.$ (9)
  3. 3.
    For the given pair $(A,B)$, the companion pair $(A_{c},B_{c})$ and the similarity $T$ are unique.
Remark 1.
Compared with the classical $\mathbb{R}/\mathbb{C}$ case, the construction is completely reversed: rather than deriving the companion realization from a characteristic/minimal polynomial computed via determinants (as usual in the classical $\mathbb{R}/\mathbb{C}$ setting), we first construct the controllable companion form directly from the controllability matrix and only then read off the companion polynomial from its last row. This route is determinant–free and sidesteps characteristic/minimal polynomials, Cayley–Hamilton, and adjugates, which do not transfer cleanly to $\mathbb{H}$.
Proof.
The proof adapts the standard companion argument for real matrices (see [2]) to the quaternionic setting.
First, note that the coefficients $a_{i}$ defined by (7) exist and are unique whenever $\mathcal{C}$ is invertible.
Now with (9),(8) and the definition of $\mathcal{C}$ we have
$\big(T^{-1}\mathcal{C}\big)_{ij}=tA^{i+j-2}B.$ (10)
Since
[…]
it follows that the matrix $T^{-1}\mathcal{C}$ is unit anti-triangular (ones on the anti-diagonal, zeros above).
The claim that $(T^{-1}AT,\;T^{-1}B)$ is in lower controllable canonical form — and hence equals $(A_{c},B_{c})$ — is proved in three steps:
Step 1 - Input vector: The first column of $\mathcal{C}$ is $B,$ hence
$T^{-1}B$ $=$ $\begin{bmatrix}tB\\ tAB\\ \vdots\\ tA^{n-1}B\end{bmatrix}=\begin{bmatrix}0\\ \vdots\\ 0\\ 1\end{bmatrix}=e_{n}=B_{c}.$
Step 2 - Superdiagonal ones: Denoting
$M$ $\coloneqq$ $T^{-1}AT,$
we prove $M$ has ones above the main diagonal.
Take the reversed controllability basis
$W$ $=$ $[A^{n-1}B,\;A^{n-2}B,\;\dots,\;B],$
which is invertible and set
$Q$ $\coloneqq$ $T^{-1}W,$
which is also invertible.
For $i=1,\dots,n-1$ and any $j=1,\dots,n$, the $(i,j)$ element of the matrix $MQ$ reads
$e_{i}^{\mathsf{T}}M\,Qe_{j}$ $=$ $e_{i}^{\mathsf{T}}T^{-1}ATT^{-1}We_{j}=e_{i}^{\mathsf{T}}T^{-1}AWe_{j}$
$=$ $e_{i}^{\mathsf{T}}T^{-1}A^{n-j+1}B=tA^{n+i-j}B,$
since $e_{i}^{\mathsf{T}}T^{-1}=tA^{i-1}$ and $AWe_{j}=A^{n-j+1}B$.
Similarly, the $(i+1,j)$ element of the matrix $Q$ reads
$e_{i+1}^{\mathsf{T}}Qe_{j}$ $=$ $e_{i+1}^{\mathsf{T}}T^{-1}We_{j}=tA^{n+i-j}B,$
since $e_{i+1}^{\mathsf{T}}T^{-1}=tA^{i}$ and $We_{j}=A^{n-j}B$.
Hence
$(e_{i}^{\mathsf{T}}M-e_{i+1}^{\mathsf{T}})\,Qe_{j}=0\textrm{ for all }j.$
Because $Q$ is invertible, we conclude $e_{i}^{\top}M=e_{i+1}^{\top}$ for $i=1,\ldots,n-1$; i.e., $M$ has ones on the superdiagonal and zeros above.
Step 3 - Last row (shift-register argument): Apply $T^{-1}$ to (7) and use $M^{k}e_{n}=T^{-1}A^{k}B$ (where $e_{n}=T^{-1}B$ commutes with $a_{k}$):
$M^{n}e_{n}\;=\;-\sum_{k=0}^{n-1}M^{k}e_{n}\,a_{k}.$ (12)
Set $v_{k}:=M^{k}e_{n}$. Since $M$ is a shift matrix, the first n1n\!-\!1 rows only shift upward; thus for each row $i$ there is a unique k=nik=n-i with $(v_{n-i})_{i}=1$ and $(v_{k})_{i}=0$ for $k\neq n-i$. Reading (12) in row $i$ gives $(v_{n})_{i}=-(v_{n-i})_{i}a_{n-i}=-a_{n-i}.$ But $(v_{n})_{i}\!=\!r_{i-1}$ (it is created by the last row via the shifts). Hence $r_{i-1}\!=\!-a_{n-i}$ for all $i$, i.e., the last row of $M$ is $a_{0},\,-a_{1},\,\dots,\,-a_{n-1}\,]$.
Therefore $M$ is the lower controllable companion form and $A_{c}:=M=T^{-1}AT,\ B_{c}:=T^{-1}B=e_{n}$.
Step 4 - Uniqueness: Since $\mathcal{C}$ is invertible, the expansion $A^{n}B=-\sum_{k=0}^{n-1}A^{k}B\,a_{k}$ determines $a_{0},\dots,a_{n-1}$ uniquely, hence $A_{c}$ is unique. With $B_{c}=e_{n}$ and ones on the superdiagonal fixed, the companion controllability matrix $\mathcal{C}_{c}=[B_{c},A_{c}B_{c},\dots,A_{c}^{n-1}B_{c}]$ is uniquely determined, and $T^{-1}\mathcal{C}=\mathcal{C}_{c}$ implies $T^{-1}=\mathcal{C}_{c}\,\mathcal{C}^{-1},$ so $T$ is unique as well. ∎
Example 1.
Let
$A$ $=$ $\begin{bmatrix}1&\mathbf{i}\\[2.0pt] \mathbf{j}&\mathbf{k}\end{bmatrix},\quad B=\begin{bmatrix}1\\[2.0pt] \mathbf{k}\end{bmatrix}$. (14)
Then the controllability matrix and its inverse are
$\mathcal{C}$ $=$ [1 1-jk-1+j] , (15)
$\mathcal{C}^{-1}$ $=$ $\tfrac{1}{4}$ [2-2k2-2k1+i+j+k-1+i-j+k]. (16)
Hence
$t=\tfrac{1}{4}\begin{bmatrix}1+\mathbf{i}+\mathbf{j}+\mathbf{k}&-1+\mathbf{i}-\mathbf{j}+\mathbf{k}\end{bmatrix}$
and
$T^{-1}$ $=$ $\tfrac{1}{4}$ [1+i+j+k-1+i-j+k2+k-2-2k], (17)
$T$ $=$ [-i-k1-i-$k$]. (18)
The lower controllable companion form is
$A_{c}$ $=$ $\begin{bmatrix}0&1\\ 1-\mathbf{i}+\mathbf{j}-\mathbf{k}&1+\mathbf{i}-\mathbf{j}+\mathbf{k}\end{bmatrix},\,B_{c}=\begin{bmatrix}0\\[2.0pt] 1\end{bmatrix}$, (19)
and the associated companion polynomial is
$a(\lambda)$ $=$ $-1+\mathbf{i}-\mathbf{j}+\mathbf{k}-(1+\mathbf{i}-\mathbf{j}+\mathbf{k})\lambda+\lambda^{2}.$ (20)
We can still have a clean annihilating identity for the lower controllable companion even though full Cayley–Hamilton fails over $\mathbb{H}.$ See [3].
Theorem 2 (Companion polynomial annihilates its companion matrix over $\mathbb{H}$).
Let $A_{c}\in\mathbb{H}^{n\times n}$ be in the companion form (4). Then the companion polynomial
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{H}[\lambda]$
annihilates $A_{c}$ from the right in the sense that
$a_{0}I+a_{1}A_{c}+\cdots+a_{n-1}A_{c}^{n-1}+A_{c}^{n}=0.$
Proof.
Set $v_{k}:=A_{c}^{\,k}e_{n}$, $k\geq 0$ and collect the first $n$ of them into the Krylov matrix
$\mathcal{K}:=[\,v_{0},\;v_{1},\;\dots,\;v_{n-1}\,]\in\mathbb{H}^{n\times n}.$
For the companion $A_{c}$, the columns $v_{0},\dots,v_{n-1}$ are linearly independent (indeed $\mathcal{K}$ is a permutation of the identity plus lower-row corrections), hence $\mathcal{K}$ is invertible.
By construction of $A_{c}$,
$A_{c}v_{k}$ $=$ $v_{k+1}\qquad(0\leq k\leq n-2),$
$A_{c}v_{n-1}$ $=$ $v_{n}=-\sum_{k=0}^{n-1}a_{k}\,v_{k},$ (21)
where the last equality is exactly the “feedback row” $a_{0},\dots,-a_{n-1}]$ applied at the wrap-around.
Left-multiplying the second relation in (21) by $A_{c}$ gives, for all $j\geq 0$
$v_{n+j}=A_{c}^{\,j}v_{n}=-\sum_{k=0}^{n-1}a_{k}\,A_{c}^{\,j}v_{k}=-\sum_{k=0}^{n-1}a_{k}\,v_{k+j}.$ (22)
Now evaluate the polynomial $a(\lambda)$ from the right on $A_{c}$ and apply it to each $v_{j}$ with $0\leq j\leq n-1$:
$\bigl(A_{c}^{n}+a_{n-1}A_{c}^{n-1}+\cdots+a_{0}I\bigr)v_{j}=v_{n+j}+\sum_{k=0}^{n-1}a_{k}v_{k+j}=0$
by (22). Thus every column of $a(A_{c})\mathcal{K}$ is zero, i.e. $a(A_{c})\mathcal{K}=0.$ Since $\mathcal{K}$ is invertible, $a(A_{c})=0$, which is the claim. ∎
Remark 2.
This is not the (generally false) quaternionic Cayley–Hamilton theorem; it is a one-sided polynomial identity tied to the companion’s shift structure.
The right spectrum of the companion matrix consists of similarity classes formed by the right zeros of the companion polynomial.
Theorem 3 (Right zeros form the right spectrum of $A_{c}$).
Let $(A_{c},B_{c})$ be in the quaternionic companion form (4) and
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{H}[\lambda]$
be the related quaternionic companion polynomial. Then
  1. 1.
    If $\lambda\in\mathbb{H}$ is a right zero of $a(\lambda),$ then $\lambda$ is also a right eigenvalue of $A_{c}$ and $v=[1,\lambda,\ldots,\lambda^{n-1}]^{\mathsf{T}}$ satisfies $A_{c}v=v\lambda$.
  2. 2.
    Consequently, $\sigma_{\mathscr{r}}(A_{c})=\bigcup_{i}[\lambda_{i}]$, where $\{\lambda_{i}\}$ is the multiset of (not necessarily distinct) right zeros of $a$.
  3. 3.
    Every left zero $\mu$ of $a(\lambda)$ (under left evaluation) is a right eigenvalue of $A_{c}$ as well.
Proof.
Step 1: Put $v\!=\!\left[1,\lambda,\ldots,\lambda^{n-1}\right]^{\mathsf{T}}$ and assume $a(\lambda)\!=\!0$. By construction,
$(A_{c}v)_{1}$ $=$ $\lambda,\ \ldots,\ (A_{c}v)_{n-1}=\lambda^{n-1},$
$(A_{c}v)_{n}$ $=$ $-(a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1})=\lambda^{n}.$
Thus $A_{c}v=v\lambda.$
Step 2: By Step (1), every right zero $\lambda_{i}$ yields a right eigenpair $A_{c}v=v\lambda_{i}$. Hence the whole similarity class $\lambda_{i}]$ lies in the right spectrum $\lambda_{i}]\subset\sigma_{\mathscr{r}}(A_{c}).$
For the converse, take a right eigenpair $A_{c}x=x\mu,$ $x\neq 0.$ From the first n1n-1 rows,
$x_{1}$ $=$ $x_{0}\mu,\ x_{2}=x_{1}\mu=x_{0}\mu^{2},\ \ldots,\ x_{n-1}=x_{0}\mu^{n-1}.$
As above $x_{0}\neq 0$. The last row gives
$-(a_{0}x_{0}+a_{1}x_{0}\mu+\cdots+a_{n-1}x_{0}\mu^{n-1})$ $=$ $x_{0}\mu^{n}.$
Right-multiply by $x_{0}^{-1}$
$-a_{0}-a_{1}(x_{0}\mu x_{0}^{-1})-\cdots-a_{n-1}(x_{0}\mu^{n-1}x_{0}^{-1})$ $=$ $x_{0}\mu^{n}x_{0}^{-1}.$
Let $\lambda:=x_{0}\mu x_{0}^{-1}.$ Then $x_{0}\mu^{k}x_{0}^{-1}=\lambda^{k}$ for all $k$, so the equality becomes
$-a_{0}-a_{1}\lambda-\cdots-a_{n-1}\lambda^{n-1}=\lambda^{n}\;\iff\;a(\lambda)=0.$
Thus every right eigenvalue $\mu$ is similar to a right zero $\lambda$ of $a(\lambda)$, i.e. $\mu\in[\lambda]$. Hence
$\sigma_{\mathscr{r}}(A_{c})$ $=$ $\textstyle{\bigcup}_{i}\,[\lambda_{i}],$
with $\{\lambda_{i}\}$ the multiset of right zeros of $a(\lambda).$
Step 3: The left zeros of $a(\lambda)$ are known to be similar to the right zeros of $a(\lambda)$ [9] and as such are also in the right spectrum of $A_{c}.$
The companion polynomial can actually be associated with any cyclic quaternionic matrix $A$ using its controllable companion form $A_{c}.$
Definition 2 (Companion polynomial of a controllable pair).
Let $(A,B)$ with $A\in\mathbb{H}^{n\times n}$ and $B\in\mathbb{H}^{n}$ satisfy that the controllability matrix $\mathcal{C}=[B,AB,\dots,A^{n-1}B]$ is invertible over $\mathbb{H}$. Let $(A_{c},B_{c})$ be the lower controllable companion form of $(A,B)$; this form is unique. The companion polynomial of $(A,B)$ is the monic quaternionic polynomial
$a(\lambda)$ $=$ $a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n},$ (23)
where $\left[-a_{0},\dots,-a_{n-1}\right]$ is the last row of $A_{c}$.
Equivalently, the coefficients $a_{0},\dots,a_{n-1}\in\mathbb{H}$ are the unique solution of
$A^{n}B\;=\;-\sum_{k=0}^{n-1}A^{k}B\,a_{k},$
which is guaranteed by the invertibility of $\mathcal{C}$.
Remark 3.
Throughout this note, we adopt the lower controllable companion form. Other orientations—upper, right, or left—are equivalent; using any of them leads to the same results (up to the obvious permutation/similarity). In particular, the companion polynomial is independent of the chosen orientation: all companion forms yield the same monic polynomial.
Theorem 4 (Right zeros form the right spectrum of $A$).
Let $(A,B)$ be a controllable quaternionic pair and $a(\lambda)$ be its quaternionic companion polynomial. Then
  1. 1.
    If $\lambda\in\mathbb{H}$ is a right zero of $a(\lambda),$ then $\lambda$ is also a right eigenvalue of $A.$
  2. 2.
    Consequently, $\sigma_{\mathscr{r}}(A)=\bigcup_{i}[\lambda_{i}]$, where $\{\lambda_{i}\}$ is the multiset of (not necessarily distinct) right zeros of $a$.
  3. 3.
    Every left zero $\mu$ of $a$ (under left evaluation) is a right eigenvalue of $A$.
Proof.
Since $A$ is similar to $A_{c}$, they share the right spectrum; the claims follow from Theorem 3. ∎
Remark 4.
Hence, the companion polynomial shares some properties of the minimal/characteristic polynomial even in quaternions, but not all! In particular, it does not annihilate $A$ in general; it only annihilates its companion form $A_{c}$. Hence $a(A_{c})=0$ but, in general, $a(A)\neq 0.$
Remark 5.
Having associated a companion polynomial to a controllable pair $(A,B)$, we may—when $A$ is cyclic—also speak of a companion polynomial of $A$. This polynomial is not intrinsic to $A$: different cyclic choices of $B$ generally produce different polynomials. Nevertheless, their right zeros agree at the level of quaternionic similarity classes: for every such $B$, the right zeros of the associated companion polynomial form the same multiset of conjugacy classes, namely the right spectrum of $A$. By contrast, once a specific pair $(A,B)$ is fixed, the lower companion realization is unique, and therefore the polynomial attached to that pair—equivalently, to its $A_{c}$—is unique.
V Control Problem: State-Feedback Pole Placement in Quaternionic Systems
We consider single-input LTI quaternionic state–space models (2) with
$\dot{x}(t)$ $=$ $Ax(t)+Bu(t),$ (24)
with $A\in\mathbb{H}^{n\times n}$ and $B\in\mathbb{H}^{n}$. We use static state feedback
$u(t)$ $=$ $-Kx(t),\qquad K\in\mathbb{H}^{1\times n},$
which yields the closed-loop dynamics
$\dot{x}(t)=A_{\mathrm{cl}}x(t),\qquad A_{\mathrm{cl}}\coloneqq A-BK.$
In the real or complex case, the method assigns $n$ (not necessarily distinct) eigenvalues of $A_{\mathrm{cl}}$. In the quaternionic setting, the situation is different: spectra must be interpreted through right eigenvalues (1) which are determined only up to quaternionic similarity (i.e., $\lambda\sim\alpha^{-1}\lambda\,\alpha$ for any $\alpha\in\mathbb{H}\setminus\{0\}$). Consequently, pole placement specifies the closed-loop right spectrum as a multiset of such similarity classes (each class being finite or infinite), or, equivalently, by choosing canonical complex representatives
$\lambda_{k}\in\mathbb{C},\quad\operatorname{Im}\lambda_{k}\geq 0,\quad k=1,\dots,n,$
counted with algebraic multiplicity. Under the usual controllability assumptions, the feedback gain $K$ can then be designed to realize the desired right-spectrum classes of $A_{\mathrm{cl}}$.
Problem formulation (Pole placement over $\mathbb{H}$).
Given $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$, and desired targets $\{\lambda_{k}\}_{k=1}^{n}\subset\mathbb{H}$ (or their canonical complex representatives), find $K\in\mathbb{H}^{1\times n}$ such that, for the closed loop $A_{\rm cl}=A-BK$,
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ $=$ $\textstyle{\bigcup}_{i}\,[\lambda_{i}],$
with similarity classes $\lambda_{k}]$ taken with the prescribed multiplicities.
Typically, the $\{\lambda_{k}\}$ are chosen as the right zeros of a desired monic closed-loop companion polynomial $a_{d}(\lambda)$. When $a_{d}\in\mathbb{R}[\lambda]$, nonreal roots appear in conjugate pairs and encode one spherical class each. For continuous time systems, pick representatives in the open left half–plane; $\operatorname{Re}$ sets decay rate and $\left|\operatorname{Im}\right|$ sets oscillation frequency.
VI Quaternionic Pole Placement via Controllable Companion Form
In the real/complex setting, state feedback is obtained by transforming a controllable pair to companion form and matching coefficients. Over $\mathbb{H}$, the same workflow applies for single-input controllable systems: use the companion polynomial—read from the last row of the companion matrix and computed over $\mathbb{H}$—and interpret poles as right-eigenvalue similarity classes. Pole placement then reduces to a coefficient update in controller coordinates, followed by mapping the gain back through the similarity transformation. Crucially, this construction avoids determinants and characteristic/minimal polynomials—as well as Cayley–Hamilton and adjugates—which do not carry over cleanly to $\mathbb{H}$.
Theorem 5 (Pole placement via controllable companion form).
Assume $(A,B)$ with $A\in\mathbb{H}^{n\times n}$, $B\in\mathbb{H}^{n}$ is controllable, and let $(A_{c},B_{c})$ be its (lower) controllable companion form via $T$ so that $A_{c}=T^{-1}AT$, $B_{c}=T^{-1}B$. Write the last row of $A_{c}$ as $\left[-a_{0},\dots,-a_{n-1}\right]$ and let $a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}$ be the companion polynomial. Given any monic
$a_{d}(\lambda)$ $=$ $d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n},$ (25)
there exists a unique feedback u=Kxu=-Kx, $K\in\mathbb{H}^{1\times n}$, such that the right spectrum of $A_{\mathrm{cl}}:=A-BK$ equals the multiset of right zeros (similarity classes, with multiplicities) of $a_{d}$. Moreover
$K$ $=$ $K_{c}\,T^{-1},$ (26)
$K_{c}$ $=$ $\big[d_{0}-a_{0}\ \ d_{1}-a_{1}\ \ \cdots\ \ d_{n-1}-a_{n-1}\big].$ (27)
Proof.
In controllable coordinates, with $K_{c}:=KT$, we have
$A_{\mathrm{cl},c}:=T^{-1}(A-BK)T=A_{c}-B_{c}K_{c}.$
Writing $K_{c}=[k_{0}\ k_{1}\ \cdots\ k_{n-1}]$, the closed-loop companion polynomial $a_{\mathrm{cl}}(\lambda)$ reads
$(a_{0}+k_{0})+(a_{1}+k_{1})\lambda+\cdots+(a_{n-1}+k_{n-1})\lambda^{n-1}+\lambda^{n}.$
Setting $k_{i}=d_{i}-a_{i}$ yields $a_{\mathrm{cl}}(\lambda)=a_{d}(\lambda)$. By the previously established spectral property of the companion form, the right zeros of $a_{\mathrm{cl}}$ are the right eigenvalues of $A_{\mathrm{cl},c}$, and therefore of $A_{\mathrm{cl}}$. Uniqueness follows since the coefficient update $k_{i}=d_{i}-a_{i}$ is bijective and $K\mapsto K_{c}=KT$ is bijective (invertible $T$). Finally $K=K_{c}T^{-1}$. ∎
Example 2 (Example 1 continued).
With the companion form (19), the similarity matrices (17), and the companion polynomial (20) in hand, we now consider several choices of desired polynomials:
(a) Real targets: $a_{d}(\lambda)=\lambda^{2}+3\lambda+2$ ($\lambda_{1}=-1,\lambda_{2}=-2$) gives
$K_{c}$ $=$ $\begin{bmatrix}3-\mathbf{i}+\mathbf{j}-\mathbf{k}&4+\mathbf{i}-\mathbf{j}+\mathbf{k}\end{bmatrix}$,
$K$ $=$ $\begin{bmatrix}2.5+\mathbf{i}+2.5\mathbf{k}&-1.5+\mathbf{i}-1.5\mathbf{k}\end{bmatrix}$, (28)
and the closed-loop matrix
$A_{\rm cl}$ $=$ $\begin{bmatrix}-1.5-\mathbf{i}-2.5\mathbf{k}&1.5+1.5\mathbf{k}\\ 2.5-2.5\mathbf{k}&-1.5-\mathbf{j}+2.5\mathbf{k}\end{bmatrix}$ (29)
has the finite right spectrum $\sigma_{\mathscr{r}}(A_{\rm cl})=\{-1,-2\}$ as desired.
(b) Complex pair: $a_{d}(\lambda)=\lambda^{2}+2\lambda+2$ ($\lambda_{1,2}=-1\pm\mathbf{i}$) yields
$K_{c}$ $=$ $\begin{bmatrix}3+\mathbf{i}-\mathbf{j}+\mathbf{k}&3-\mathbf{i}+\mathbf{j}-\mathbf{k}\end{bmatrix}$,
$K$ $=$ $\begin{bmatrix}2+\mathbf{i}+2\mathbf{k}&-1+\mathbf{i}-\mathbf{k}\end{bmatrix}$, (30)
and
$A_{\rm cl}$ $=$ $\begin{bmatrix}-1-\mathbf{i}-2\mathbf{k}&1+\mathbf{k}\\ 2-2\mathbf{k}&-1-\mathbf{j}+2\mathbf{k}\end{bmatrix}$. (31)
Here $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$; equivalently, the entire spherical (similarity) class $\left[-1+\mathbf{i}\right]$ is assigned as expected.
(c) Quaternionic coefficients: Demanding specific quaternion poles (e.g., $\lambda_{1}=-1+\mathbf{j}$ and $\lambda_{2}=-2+\mathbf{k}$) sets (rounded)
$a_{d}(\lambda)$ $=$ $2.7-\mathbf{i}-1.3\mathbf{j}+0.33\mathbf{k}$ (32)
$+(3-0.67\mathbf{i}-0.33\mathbf{j}-0.33\mathbf{k})\lambda+\lambda^{2}.$
Here $a_{d}(\lambda)\in\mathbb{H}[\lambda]$ has nonreal (noncentral) coefficients; evaluating $a_{d}(A)$ is then quaternionic, and we obtain
$K_{c}$ $=$ ${\begin{bmatrix}3.7\!-\!2\mathbf{i}\!-\!0.33\mathbf{j}\!-\!0.67\mathbf{k},&4\!+\!0.33\mathbf{i}\!-\!1.3\mathbf{j}\!+\!0.67\mathbf{k}\end{bmatrix}}$,
$K$ $=$ ${\begin{bmatrix}3.3\!+\!0.33\mathbf{j}\!+\!2.7\mathbf{k},&-2\!+\!1.7\mathbf{i}\!+\!0.33\mathbf{j}\!-\!0.67\mathbf{k}\end{bmatrix}}$,
and
$A_{\rm cl}$ $=$ ${\begin{bmatrix}-2.3\!-\!0.33\mathbf{j}\!-\!2.7\mathbf{k},&2\!-\!0.67\mathbf{i}\!-\!0.33\mathbf{j}\!+\!0.67\mathbf{k}\\ 2.7\!+\!0.33\mathbf{i}\!+\!\mathbf{j}\!-\!3.3\mathbf{k},&\!-\!0.67\!+\!0.33\mathbf{i}\!-\!1.7\mathbf{j}\!+\!3\mathbf{k}\end{bmatrix}}$.
Its right spectrum is $\sigma_{\mathscr{r}}(A_{\mathrm{cl}})=\left[-1+\mathbf{i}\right]\cup[-2+\mathbf{i}]=\left[-1+\mathbf{j}\right]\cup\left[-2+\mathbf{k}\right]$ so $1\!+\!\mathbf{j},\,-2\!+\!\mathbf{k}\in\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ confirming that the controller-form coefficient update achieves the prescribed classes even with quaternionic coefficients.
VII Quaternionic Ackermann Formula
In this section, we show that the Ackermann formula carries over to quaternionic systems. Unlike the companion-form coefficient-update method, it is not fully general: it applies when the desired closed-loop companion polynomial has real coefficients. Practically, this means we assign right eigenvalue classes (spherical similarity classes) with chosen multiplicities—still broad enough for design, but not literally unrestricted. Algorithmically, the recipe is identical to the classical case (same controllability matrix selection and one-shot gain computation). At the same time, the proof is different in substance: over $\mathbb{H}$ we cannot appeal to the determinant or Cayley–Hamilton, so we justify the formula via the controllable companion relations and a central (real-coefficient) polynomial calculus.
Theorem 6 (Quaternionic Ackermann formula).
Let $A\in\mathbb{H}^{n\times n}$ and $B\in\mathbb{H}^{n}$ be such that the pair $(A,B)$ is controllable with $\mathcal{C}=\left[B,AB,\ldots,A^{n-1}B\right]\in\mathbb{H}^{n\times n}$. Let
$a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n}\in\mathbb{R}[\lambda],$ (33)
be the desired monic closed-loop companion polynomial with real coefficients and let $e_{n}=[0,\dots,0,1]^{\mathsf{T}}\in\mathbb{R}^{n}$ be the last canonical basis vector. Then the state-feedback row vector
$K$ $=$ $e_{n}^{\mathsf{T}}\mathcal{C}^{-1}a_{d}(A),$ (34)
with u=Kxu=-Kx yields the closed-loop matrix ABKA-BK whose right spectrum equals the similarity classes induced by the right roots of $a_{d}(\lambda)$
$\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$ $=$ $\bigcup_{i=1}^{n}[\lambda_{i}],$ (35)
where $\{\lambda_{i}\}_{i=1}^{n}$ are the right roots of $a_{d}(\lambda)$ (listed with multiplicity).
Remark 6.
The Ackermann step is valid only when $a_{d}\in\mathbb{R}[\lambda]$, because real scalars lie in the center of $\mathbb{H}$, ensuring that $a_{d}(A)$ is unambiguous. In this case: closed-loop poles are real or occur as complex-conjugate pairs; left/right zeros coincide; and any nonreal root $\gamma\in\mathbb{C}$ actually stands for the entire spherical class $\gamma]\subset\mathbb{H}$ (choosing $\gamma,\bar{\gamma}$ assigns that class).
For an $n$th-order single-input design, each real class uses one degree and each nonreal class uses two; with $r$ real and $s$ nonreal classes, $2s+r\leq n.$
The following Lemma is important in the Ackermann formula proof.
Lemma 1 (Polynomial Intertwining).
Let $p(\lambda)\in\mathbb{R}[\lambda]$ be a polynomial with real coefficients. If an invertible matrix $T$ satisfies
$AT$ $=$ $T\!A_{c}$ (36)
then
$p\!\left(A\right)\!T$ $=$ $Tp\!\left(A_{c}\right).$ (37)
Proof.
From $AT=TA_{c}$ it follows by induction that
$A^{k}T$ $=$ $TA_{c}^{\,k}\quad(k=0,1,\dots).$ (38)
Write $p(\lambda)=\sum_{k=0}^{n}p_{k}\lambda^{k}$ with $p_{k}\in\mathbb{R}.$ Then
$p(A)T$ $=$ $\sum_{k=0}^{n}p_{k}A^{k}T=\sum_{k=0}^{n}p_{k}T\!A_{c}^{\,k}$
$=$ $T\sum_{k=0}^{n}p_{k}A_{c}^{\,k}=Tp(A_{c}),$
where we used that real coefficients lie in the center $Z(\mathbb{H})=\mathbb{R}$ and commute with $T$ and with powers of $A_{c}.$
Remark 7.
The lemma generally fails if $p\notin\mathbb{R}[\lambda]$. Although one may of course define $p(A_{c})=p(T^{-1}AT)$ by direct substitution, the intertwining identity
$p(A)\,T$ $\stackrel{{\scriptstyle?}}{{=}}$ $T\,p(A_{c}).$ (39)
need not hold once any coefficient of $p$ is a nonreal quaternion. Indeed, take, $A=\mathbf{i}$ and $T=\mathbf{k}$ so that $A_{c}=T^{-1}AT=-\mathbf{i},$ and $p(\lambda)=\lambda+\mathbf{i}$. We have $p(A)T=(\mathbf{i}+\mathbf{i})k=2\mathbf{i}\mathbf{k}=-2\mathbf{j}$ but $Tp(A_{c})=k(-\mathbf{i}+\mathbf{i})=0$, hence $p(A)T\neq Tp(A_{c})$. This illustrates why the Ackermann proof below requires $p\in\mathbb{R}[\lambda]$: real coefficients ensure unambiguous evaluation and permit the coefficient–matrix “swap” used in the intertwining step.
Proof of Theorem 6.
Classical proofs over $\mathbb{R}$ or $\mathbb{C}$ lean on determinants and adjoints [11, 4], characteristic polynomials [4, 8], or Cayley–Hamilton—steps [11, 4]; see also Ackermann’s original paper [1]. These steps do not transfer verbatim to $\mathbb{H}$ because of noncommutativity and left/right evaluation. We therefore avoid those tools and proceed differently:
Step 1: Controllable form and coefficient matching. Let an invertible similarity $T$ put $(A,B)$ into lower controllable companion form:
$A_{c}$ $=$ $T^{-1}AT,\qquad B_{c}=T^{-1}B.$
Write the open-loop companion polynomial as
$a(\lambda)=a_{0}+a_{1}\lambda+\cdots+a_{n-1}\lambda^{n-1}+\lambda^{n}.$
Given the desired (real) closed-loop polynomial
$a_{d}(\lambda)=d_{0}+d_{1}\lambda+\cdots+d_{n-1}\lambda^{n-1}+\lambda^{n},$ (40)
the controllable-form feedback row
$K_{c}$ $=$ $\begin{bmatrix}d_{0}-a_{0}&\ldots&d_{n-1}-a_{n-1}\end{bmatrix}$ (41)
assigns the right eigenvalues of $A_{c}-B_{c}K_{c}$ to the roots of $a_{d}$ (counted with multiplicity, interpreted as right-eigenvalue classes). The feedback in the original coordinates is
$K=K_{c}T^{-1}$
Step 2: Controllable-form Ackermann equals coefficient matching. Let $\mathcal{C}_{c}=\left[B_{c},A_{c}B_{c},\ldots,A_{c}^{n-1}B_{c}\right]$ be the controllability matrix of the controllable form (4), and $e_{n}$ the last canonical basis vector. The controllable-form Ackermann expression
$K_{c}$ $=$ $e_{n}^{\mathsf{T}}\mathcal{C}_{c}^{-1}a_{d}(A_{c})$ (42)
produces exactly the same $K_{c}$ as above. The special structure of $\mathcal{C}_{c}^{-1}$ in the controllable companion basis implies $e_{n}^{\top}\mathcal{C}_{c}^{-1}=e_{1}^{\top}$. Using this in (42), we have
$K_{c}$ $=$ $e_{1}^{\mathsf{T}}a_{d}(A_{c}).$
Furthermore, the shape of $A_{c}$ gives
$e_{1}^{\mathsf{T}}A_{c}^{k}$ $=$ $e_{k+1}^{\top}\quad(k=0,\dots,n-2),$
$e_{1}^{\mathsf{T}}A_{c}^{n}$ $=$ $\big[-a_{0},\dots,-a_{n-1}\big],$
so, since $e_{1}$ commutes with the $d_{i}$ (all real),
$K_{c}$ $=$ $e_{1}^{\top}a_{d}(A_{c})$
$=$ $\big[d_{0},\dots,d_{n-1}\big]+\big[-a_{0},\dots,-a_{n-1}\big]$
$=$ $\begin{bmatrix}d_{0}-a_{0}&\ldots&d_{n-1}-a_{n-1}\end{bmatrix}$
which is exactly the coefficient-matching gain (41).
Step 3 — Coordinate-free form and mapping back: Let $\mathcal{C}$ be the controllability matrix (3) of $(A,B)$. From the similarity, for $k\geq 0$ we have
$B$ $=$ $TB_{c},\quad A^{k}T=TA_{c}^{\,k}.$ (43)
Hence
$\mathcal{C}=T\mathcal{C}_{c}\implies\mathcal{C}^{-1}=\mathcal{C}_{c}^{-1}T^{-1}\;\mathrm{and\;}\;T=\mathcal{C}\,\mathcal{C}_{c}^{-1}.$ (44)
Because $a_{d}\in\mathbb{R}[\lambda],$ the Lemma 1 yields
$T^{-1}\,a_{d}(A)$ $=$ $a_{d}(A_{c})\,T^{-1}.$ (45)
Therefore
$K$ $=$ $e_{n}^{\mathsf{T}}\mathcal{C}^{-1}a_{d}(A)=e_{n}^{\top}C_{c}^{-1}\,T^{-1}a_{d}(A)$
$=$ $\big(e_{n}^{\top}C_{c}^{-1}a_{d}(A_{c})\big)\,T^{-1}=K_{c}T^{-1}.$
since $K_{c}=e_{n}^{\mathsf{T}}C_{c}^{-1}a_{d}(A_{c}).$
Thus, the basis-invariant identity (VII) produces exactly the same feedback as the controllable-form construction, completing the proof. ∎
Example 3 (Example 2 continued).
Continuing the previous example, we already have the controllability inverse $\mathcal{C}^{-1}$ from (16); we now consider specific target polynomials.
(a) Real targets: $a_{d}(\lambda)=\lambda^{2}+3\lambda+2$ ($\lambda_{1}=-1,\lambda_{2}=-2$) gives
$a_{d}(A)=\begin{bmatrix}6+\mathbf{k}&4\mathbf{i}-\mathbf{j}\\ -\mathbf{i}+4\mathbf{j}&1+2\mathbf{k}\end{bmatrix}$.
and the Ackermann formula (34) yields
$K$ $=$ $\begin{bmatrix}2.5+\mathbf{i}+2.5\mathbf{k}&-1.5+\mathbf{i}-1.5\mathbf{k}\end{bmatrix}$,
which equals (28). The resulting closed-loop matrix $A_{\rm cl}=A-BK$ is identical to (29) and its right spectrum is $\sigma_{\mathscr{r}}(A_{\rm cl})=\{-1,-2\}$ as desired.
(b) Complex pair: $a_{d}(\lambda)=\lambda^{2}+2\lambda+2$ ($\lambda_{1,2}=-1\pm\mathbf{i}$) gives rise to
$a_{d}(A)$ $=$ $\begin{bmatrix}5+\mathbf{k}&3\mathbf{i}-\mathbf{j}\\ -\mathbf{i}+3\mathbf{j}&1+\mathbf{k}\end{bmatrix}$,
and
$K=\begin{bmatrix}2+\mathbf{i}+2\mathbf{k}&-1+\mathbf{i}-\mathbf{k}\end{bmatrix}$
which equals (30). The closed-loop matrix $A_{\rm cl}$ is identical to (31) and $\{-1\!+\!\mathbf{i},-1\!-\!\mathbf{i}\}\subset\sigma_{\mathscr{r}}(A_{\mathrm{cl}})$.
(c) Quaternionic coefficients, for which the formula fails: Demanding quaternion poles $\lambda_{1}=-1+\mathbf{j}$, $\lambda_{2}=-2+\mathbf{k}$ and using $a_{d}(\lambda)$ as in (32), the Ackermann returns
$K$ $=$ ${\begin{bmatrix}3.5\!+\!1.3\mathbf{i}\!-\!0.5\mathbf{j}\!+\!3\mathbf{k},&-1.5\!+\!2\mathbf{i}\!-\!0.5\mathbf{j}\!-\!1.7\mathbf{k}\end{bmatrix}}$.
However, this feedback results in
$A_{\mathrm{cl}}$ $=$ ${\begin{bmatrix}-2.5\!-\!1.3\mathbf{i}\!+\!0.5\mathbf{j}\!-\!3\mathbf{k},&1.5\!-\!\mathbf{i}\!+\!0.5\mathbf{j}\!+\!1.7\mathbf{k}\\[2.0pt] 3\!-\!0.5\mathbf{i}\!-\!0.33\mathbf{j}\!-\!3.5\mathbf{k},&1.7\!-\!0.5\mathbf{i}\!-\!2\mathbf{j}\!+\!2.5\mathbf{k}\end{bmatrix}}$
whose right spectrum is $0.48+0.85\mathbf{i}\,]\cup[\,-3.7+2.4\mathbf{i}\,]$, confirming that when $a_{d}\notin\mathbb{R}[\lambda]$, the Ackermann construction does not enforce the prescribed classes.
VIII Conclusion
This note developed a quaternionic framework for state-feedback pole placement. We (i) formalized controllable companion forms over $\mathbb{H}$ and their companion polynomials, which right-annihilate the companion matrix, and (ii) gave two practical designs: coefficient matching in controllable coordinates and a coordinate-free Ackermann construction valid for real (central) targets. The method keeps the convenience of field-based workflows while addressing quaternionic specifics—companion polynomial instead of characteristic/minimal, right-eigenvalue similarity classes, and order-sensitive polynomial evaluation.
Although we focused on single-input systems, extension to multi-input appears straightforward via block companion/Brunovsky forms and the same central-polynomial calculus—mainly algebraic bookkeeping. Future work will focus on observer duals, robustness, the numerical conditioning of quaternionic companion computations, and output-feedback designs.